OpenStudy (anonymous):

Claire has received scores of 85, 88, 87, and 90 on her algebra tests. What is the minimum score she must receive on the fifth test to have an overall test score average of at least 87? (Hint: The average of a list of numbers is their sum divided by the number of numbers in the list.)

OpenStudy (anonymous):

(85+88+87+90+x)/5=87 (350+x)/5=87 350+x=435 x=85

OpenStudy (anonymous):

Alright thanks for the help